# Robustness of Dynamic Quantum Control: Differential Sensitivity Bounds

### v1

@@ IV.3 Effect of Control Type and Target Gate
Fig. 5a shows robustness versus fidelity error comparison of controllers for Problems $4$ and $7$. While both systems are five-qubit registers with Ising coupling between adjacent qubits, in Problem $4$ each qubit is individually controllable while in Problem $7$ the control fields act globally, i.e., the controls affect simultaneous rotations on all qubits but a linear detuning is applied to the qubit register. Though the control implementation for Problem $7$ might seem less effective due to the reduced degrees of freedom, as Fig. 5a shows, this is not necessarily the case. While for Problem $4$ controllers with $t_{f}=15$ exhibit the lowest bounds $\mathbf{B_{vu}}$ on the differential sensitivity (see Fig. 5c), the Problem $7$ controllers (see Fig. 5b) have similar or better fidelity, with error on the order of $10^{-5}$ or less, for only slighter greater differential sensitivity. As with the previous case, this suggests that optimizing for a control paradigm with reduced degrees of freedom has the potential to yield controllers that exhibit good performance and low sensitivity with the same benefits in reduction of computational overhead and increased robustness in implementation.
We also examine the effect of differing target gates on systems of the same size with the same controller implementation. Fig. 6 shows the plots of $\mathbf{B_{vu}}$ versus $\varepsilon$ for the three-qubit problems, grouped by control implementation. Fig. 6a shows that for the QFT target gate. Controllers for the Heisenberg-coupled system of Problem $5$ perform better than controllers for the Ising-coupled system of Problem $2$. However, the smallest bounds on the differential sensitivity are achieved for the Ising system. A comparison of the plots for Problems $5$ and $6$ shows that the effect of the target gate on the resulting robustness and performance of the controllers is negligible with both controllers clustering in the same area of the plot. Fig. 6b shows the same for limited control implementation of Problems $8$ and $9$. The same trend is obtained in this plot, so, at least for the problems considered here, the effect of the target gate has no noticeable impact on the performance-robustness characteristics. Rather the $t_{f}/\kappa$ ratio has a much more noticeable impact on the production of controllers with desirable properties versus those on the other end of the spectrum.
@@ V Log-Sensitivity and Fundamental Limitations
(a) Plot of $\|s\|$, $\mathbf{B_{vu}}$, and $\varepsilon$ versus controller index on a semi-log scale for the three-chain Problem $9$, $t_{f}=10$, $\kappa=64$.
(b) Plot of $\|s\|$, $\mathbf{B_{vu}}$, and $\varepsilon$ versus controller index on a semi-log scale for the four-chain Problem $4$, $t_{f}=12$, $\kappa=40$.
@@
Figure 7: Plot of log-sensitivity, sensitivity bounds, and nominal fidelity error versus controller index. The top plot shows the trade-off between log-sensitivity and fidelity error expected from classical control, while the bottom plot is one of the two exceptions in the data set.
@@
The previous sections have shown that there is generally no trade-off between the differential sensitivity bounds and the performance as measured by the fidelity error. For sake of comparison with classical control limitations, we analyze the trade-off (or lack thereof) between performance (as measured by $\varepsilon$) and robustness (as measured by the log-sensitivity), proceeding along the same lines as that done in Refs. O’Neil et al., 2023b, c. The log-sensitivity is essentially the differential sensitivity of the logarithm of the error, which equates to the differential sensitivity of the error divided by the error. Note that the definitions in Refs. O’Neil et al., 2023b and O’Neil et al., 2023c differ slightly. The former definition aligns more closely with classical control theory but fails for perturbations around a parameter with nominal value 00. To circumvent this problem the definition of the log-sensitivity for a perturbation structured as $\hat{H}_{\mu}$ was amended in Ref. O’Neil et al., 2023c. This is the definition chosen here:
@@
$s_{\mu}=\left.\frac{\partial\ln(\tilde{\varepsilon}_{\mu})}{\partial\delta}\right|_{\delta=0}=\frac{1}{\varepsilon}\zeta_{\mu}.$ (22)
@@
To streamline the analysis, we take the $2$-norm of the vector of $M+1$ possible $s_{\mu}$ for a given controller as $\|s\|=\sqrt{\sum_{\mu=0}^{M}s_{\mu}}$. This provides a single value to use as measure of the log-sensitivity for a given controller. We then test the level of concordance between the error and $\|s\|$ based on the Kendall $\tau$ rank correlation coefficient. Specifically, we execute a one-tailed test for anti-concordance with a null hypothesis $\mathcal{H}_{0}$ of no correlation between $\varepsilon$ and $\|s\|$ and an alternative hypothesis $\mathcal{H}_{-}$ of negative rank correlation. We test for concordance in the same manner, but with an alternative hypothesis $\mathcal{H}_{+}$ indicated by positive rank correlation. For both tests we set the significance level at $95\%$ ($p=0.05$).
@@
Of the $86$ total test cases, $11$ fail to meet the significance threshold. Of the $75$ remaining tests, all but two reject $\mathcal{H}_{0}$ in favor of $\mathcal{H}_{-}$. The two cases that reject $\mathcal{H}_{0}$ in favor of $\mathcal{H}_{+}$ and indicate a non-classical trend are the Problem $3$ and $t_{f}=12$ cases. Fig. 7a shows the trend between the log-sensitivity, as measured by $\|s\|$, and the differential sensitivity, as measured by the bound $\mathbf{B_{vu}}$, with the fidelity error for Problem $9$, $t_{f}=10$, and $\kappa=64$. The negative trend between the log-sensitivity and fidelity error is readily apparent and borne out the by the Kendall $\tau$ of $-0.693$. Conversely, Fig. 7b shows one of the two non-classical cases for Problem $3$, $t_{f}=12$, and $\kappa=40$. The positive correlation between $\|s\|$ and $\varepsilon$ is apparent between controller index $20$ though $80$.
@@
In summary, despite the positive trend between the differential sensitivity and $\varepsilon(t_{f})$, the expected classical limitations between the performance and robustness, as measured by the log-sensitivity observed in earlier work still hold in the main. Future work should consider which metric, the pure differential sensitivity or the normalized log-sensitivity, is a more useful measure of robustness in the context of quantum control.
@@ VI Conclusion
There are limitations to the applicability of robustness measures based on the differential sensitivity. It is likely to be a useful measure to assess and compare robustness of controllers in the high-fidelity, small perturbation regime, such as the implementation of high-fidelity quantum gates subject to small uncertainties. For problem involving large perturbations and larger errors, other robustness-infidelity measures may be useful to assess robustness post-synthesis Khalid et al. (2023a). More work is necessary to compare different robustness measures to establish when they should be used and under which conditions optimization for robustness and fidelity is beneficial.
Furthermore, we only considered structured perturbations to the system or control Hamiltonians, which are a significant source of error in many applications, but there are other sources of error, which have not been considered here. Some, such as decoherence, can be modeled using structured perturbations O’Neil et al. (2023c). Other errors such as those arising from imperfect realization of piecewise-constant controls, e.g., due to rise and settling times or bandwidth limitations may be less amenable to treatment using structured perturbations and should be considered independently using complementary methods, e.g., to compensate for nonlinear distortions in the controls due to hardware limitations Singh et al. (2023).
@@ Acknowledgements
We acknowledge the support of the Supercomputing Wales project (SCW2103), which is part-funded by the European Regional Development Fund (ERDF) via the Welsh Government.
@@
The positive correlation between the upper bounds on the differential sensitivity and the nominal fidelity error $\varepsilon(t_{f})$ calls into question whether the systems under consideration circumvent the fundamental limitations on performance and sensitivity established by classical control Safonov et al. (1981). Specifically, in a classical frequency domain context, there is a trade-off between the tracking error and the normalized logarithmic sensitivity of the tracking error, in the form of the identity $S(s)+T(s)=I$, where $S(s)$ is proportional to the tracking error and $T(s)$ is a measure of the normalized logarithmic-sensitivity (Eq. (22)) of the closed loop system to parameter variation Dorf (2011). For a purely sinusoidal input, $s=i\omega$ and $S(i\omega)$ and $T(i\omega)$ are the frequency-dependent gains from the reference input to the error signal and system output, respectively, obtained via the Laplace transform.

### v2

@@ IV.3 Effect of Control Type and Target Gate
Fig. 5a shows robustness versus fidelity error comparison of controllers for Problems $4$ and $7$. While both systems are five-qubit registers with Ising coupling between adjacent qubits, in Problem $4$ each qubit is individually controllable while in Problem $7$ the control fields act globally, i.e., the controls affect simultaneous rotations on all qubits but a linear detuning is applied to the qubit register. Though the control implementation for Problem $7$ might seem less effective due to the reduced degrees of freedom, as Fig. 5a shows, this is not necessarily the case. While for Problem $4$ controllers with $t_{f}=15$ exhibit the lowest bounds $\mathcal{B}_{\text{vu}}$ on the differential sensitivity (see Fig. 5c), the Problem $7$ controllers (see Fig. 5b) have similar or better fidelity, with error on the order of $10^{-5}$ or less, for only slighter greater differential sensitivity. As with the previous case, this suggests that optimizing for a control paradigm with reduced degrees of freedom has the potential to yield controllers that exhibit good performance and low sensitivity with the same benefits in reduction of computational overhead and increased robustness in implementation.
We also examine the effect of differing target gates on systems of the same size with the same controller implementation. Fig. 6 shows the plots of $\mathcal{B}_{\text{vu}}$ versus $\varepsilon$ for the three-qubit problems, grouped by control implementation. Fig. 6a shows that for the QFT target gate. Controllers for the Heisenberg-coupled system of Problem $5$ perform better than controllers for the Ising-coupled system of Problem $2$. However, the smallest bounds on the differential sensitivity are achieved for the Ising system. A comparison of the plots for Problems $5$ and $6$ shows that the effect of the target gate on the resulting robustness and performance of the controllers is negligible with both controllers clustering in the same area of the plot. Fig. 6b shows the same for limited control implementation of Problems $8$ and $9$. The same trend is obtained in this plot, so, at least for the problems considered here, the effect of the target gate has no noticeable impact on the performance-robustness characteristics. Rather the $t_{f}/\kappa$ ratio has a much more noticeable impact on the production of controllers with desirable properties versus those on the other end of the spectrum.
@@ V Log-Sensitivity and Fundamental Limitations
Figure 7: Plot of log-sensitivity, sensitivity bounds, and nominal fidelity error versus controller index. (a) Plot of $\|\mathcal{S}\|$, $\mathcal{B}_{\text{vu}}$, and $\varepsilon$ versus controller index on a semi-log scale for the three-chain Problem $9$, $t_{f}=10$, $\kappa=64$. (b) Plot of $\|\mathcal{S}\|$, $\mathcal{B}_{\text{vu}}$, and $\varepsilon$ versus controller index on a semi-log scale for the four-chain Problem $3$, $t_{f}=12$, $\kappa=40$. The plot in (a) shows the trade-off between log-sensitivity and fidelity error expected from classical control, while (b) shows one of the two exceptions in the data set.
@@
The previous sections have shown that there is generally no trade-off between the differential sensitivity bounds and the performance as measured by the fidelity error. For sake of comparison with classical control limitations, we analyze the trade-off (or lack thereof) between performance (as measured by $\varepsilon$) and robustness (as measured by the log-sensitivity), proceeding along the same lines as that done in Refs. 52, 53. The log-sensitivity is essentially the differential sensitivity of the logarithm of the error, which equates to the differential sensitivity of the error divided by the error. Note that the definitions in Refs. 52 and 53 differ slightly. The former definition aligns more closely with classical control theory but fails for perturbations around a parameter with nominal value 00. To circumvent this problem, the definition of the log-sensitivity for a perturbation structured as $H_{\mu}$ was amended in Ref. 53. This is the definition used here:
@@
$\mathcal{S}_{\mu}=\left.\frac{\partial\ln(\tilde{\varepsilon}_{\mu})}{\partial\delta}\right|_{\delta=0}=\frac{1}{\varepsilon}\zeta_{\mu}.$ (21)
@@
Now let the unit vector in the parameter space $\mathbf{s}_{\mu}$ only take values in $\set{e_{k}}_{m=0}^{M}$ where ${e_{k}}$ is a natural basis vector (i.e. $\mathbf{s}_{1}=(1,\ldots,0)^{T}$ through $\mathbf{s}_{M}=(0,\cdots,1)^{T}$). Then $\mathcal{S}_{\mu}$ is the log-sensitivity of the fidelity error to uncertainty in one of the principal directions $\set{0,1,\ldots,M}$. To make the analysis tractable, we take the $2$-norm of the vector of $M+1$ possible log-sensitivity values for a given controller as $\|\mathcal{S}\|=\sqrt{\sum_{\mu=0}^{M}\mathcal{S}^{2}_{\mu}}$. This provides a single value to use as measure of the log-sensitivity for a given controller. We then test the level of concordance between the error and $\|\mathcal{S}\|$ based on the Kendall $\tau$ rank correlation coefficient. Specifically, we execute a one-tailed test for anti-concordance with a null hypothesis $\mathcal{H}_{0}$ of no correlation between $\varepsilon$ and $\|\mathcal{S}\|$ and an alternative hypothesis $\mathcal{H}_{-}$ of negative rank correlation. For the opposite conclusion, we test for concordance in the same manner, but with an alternative hypothesis $\mathcal{H}_{+}$ indicated by positive rank correlation. For both tests we set the significance level at $95\%$ ($p=0.05$).
@@
Of the $86$ total test cases, $11$ fail to meet the significance threshold. Of the $75$ remaining tests, all but two reject $\mathcal{H}_{0}$ in favor of $\mathcal{H}_{-}$. The two cases that reject $\mathcal{H}_{0}$ in favor of $\mathcal{H}_{+}$ and indicate a non-classical trend are the Problem $3$ and $t_{f}=12$ cases. Fig. 7a shows the trend between the log-sensitivity, as measured by $\|\mathcal{S}\|$, and the differential sensitivity, as measured by the bound $\mathcal{B}_{\text{vu}}$, with the fidelity error for Problem $9$, $t_{f}=10$, and $\kappa=64$. The negative trend between the log-sensitivity $\|\mathcal{S}\|$ and fidelity error $\varepsilon$ is readily apparent and borne out the by the Kendall $\tau$ of $-0.693$. Conversely, Fig. 7b shows one of the two non-classical cases for Problem $3$, $t_{f}=12$, and $\kappa=40$. The positive correlation between $\|\mathcal{S}\|$ and $\varepsilon$ is not readily apparent visually but is borne out by a Kendall $\tau$ of $0.270$.
@@
In summary, despite the positive trend between the differential sensitivity and $\varepsilon(t_{f})$, the expected classical limitations between the performance and robustness, as measured by the log-sensitivity observed in earlier work [52],[53] still hold in the main. Future work should consider which metric, the pure differential sensitivity or the normalized log-sensitivity, is a more useful measure of robustness in the context of quantum control.
@@ VI Conclusion
There are limitations to the applicability of robustness measures based on the differential sensitivity. It is likely to be a useful measure to assess and compare robustness of controllers in the high-fidelity, small perturbation regime, such as the implementation of high-fidelity quantum gates subject to small uncertainties. For problem involving large perturbations and larger errors, other robustness-infidelity measures may be useful to assess robustness post-synthesis [16]. More work is necessary to compare different robustness measures to establish when they should be used and under which conditions optimization for robustness and fidelity is beneficial.
Furthermore, we only considered structured perturbations to the system or control Hamiltonians, which are a significant source of error in many applications, but there are other sources of error, which have not been considered here. Some, such as decoherence, can be modeled using structured perturbations [53]. Other errors such as those arising from imperfect realization of piecewise-constant controls, e.g., due to rise and settling times or bandwidth limitations may be less amenable to treatment using structured perturbations and should be considered independently using complementary methods, e.g., to compensate for nonlinear distortions in the controls due to hardware limitations [54].
Data Availability
All controller data, robustness analysis code, and results data associated with this manuscript is available at https://qyber.black/spinnet/code-differential-sensitivity-bounds-for-dynamic-control.
@@ Acknowledgements
We acknowledge the support of the Supercomputing Wales project (SCW2103), which is part-funded by the European Regional Development Fund (ERDF) via the Welsh Government. C.A.W., F.C.L. and S.M.S. acknowledge funding from EPSRC via grant number EP/Y004728/1.
Author Declarations
The authors have no conflicts to disclose.
@@
The positive correlation between the upper bounds on the differential sensitivity and the nominal fidelity error $\varepsilon(t_{f})$ calls into question whether the systems under consideration circumvent the fundamental limitations on performance and sensitivity established by classical control [55]. Specifically, in a classical frequency domain context, there is a trade-off between the tracking error and the normalized logarithmic sensitivity of the tracking error, in the form of the identity $S(s)+T(s)=I$, where $S(s)$ is proportional to the tracking error and $T(s)$ is a measure of the normalized logarithmic-sensitivity (Eq. (21)) of the closed loop system to parameter variation [56]. For a purely sinusoidal input, $s=i\omega$ and $S(i\omega)$ and $T(i\omega)$ are the frequency-dependent gains from the reference input to the error signal and system output, respectively, obtained via the Laplace transform.